Why does reverse saturation current increase with temperature? How fast does it increase for silicon?
1
Expert's answer
2020-05-04T12:49:58-0400

The return current is significantly dependent on temperature, while the relative change in the forward current with temperature is negligible. With increasing temperature, the barrier height qφc decreases and the concentration of minority charge carriers increases exponentially, as a result of which the saturation current increases with increasing temperature. An increase in the saturation current with increasing temperature is practically the main factor determining the temperature limit of operation of rectifier diodes. At a certain temperature, the intrinsic conductivity of the diode material becomes comparable with the impurity, while the number of minority charge carriers is comparable to the number of main ones and the rectification sharply worsens. Therefore, the upper limit temperature Tmax is higher for diodes made of semiconductor materials with a wider band gap. For example, for germanium (ΔE = 0.72 eV) T= 348-363 K, for silicon (ΔE = 1.12 eV) T = 423-473 K.